OpenStudy (anonymous):

solve for d: b= (8d- c)/d

OpenStudy (campbell_st):

well split the fraction on the right hand side \[b = \frac{8d}{d} - \frac{c}{d} \] which gives \[b = 8 - \frac{c}{d} \] then \[b - 8 = -\frac{c}{d} \] I'll leave you to finish the problem... the initial step was to split and simplify the fraction. I hope it makes sense